Rob and his brother mixed together two types of soil to make 15 m3 of soil with a 26% silt content. They used 8 m3 of a soil with 12% silt content and 7m3 of another type of soil. What was the silt content of the second type of soil?
Answer by checkley77(12844)

.12*8+7x=.26*15
.96+7x=3.9
7x=3.9-.96
7x=2.94
x=2.94/7
x=.42% silt for the 7m3 soil.
Proof:
.12*8+7*.42=.26*15
.96+2.94=3.9
3.9=3.9